# Beyond the Golden Data: Resolving the Motion-Vision Quality Dilemma via Timestep Selective Training

### 26 Mar 2026 v1

@@ 1 Introduction
In this work, we take a different perspective—shifting the question from which videos should be kept to how can we use imperfect videos more effectively? We begin by analyzing the learning behavior of video diffusion models and make two key observations. First, the denoising process exhibits hierarchical learning: motion emerges at high-noise timesteps while details refine at low-noise timesteps. Second, and more critically, through gradient-based analysis we discover that quality-imbalanced samples can match golden data’s gradients at appropriate timesteps. This reveals that imbalanced data need not be discarded but should be strategically allocated to matching denoising phases where they provide equivalent learning signals. Building upon this observation, we propose timestep selection in the training process and instantiate it through Timestep-aware Quality Decoupling. Instead of sampling timesteps equally across all data, we adjust the sampling distribution based on each video’s quality profile. High-MQ videos are biased toward higher-noise timesteps, whereas high-VQ videos are biased toward lower-noise timesteps. This design allows the model to leverage imperfect data effectively, aligning each sample with the phase of diffusion training where it provides the most benefit.
@@ 2.2 Data Selection and Curation
Data curation plays an important role in large-scale generative modeling. In vision-language modeling, both curated datasets [35] and recent model-driven selection approaches [11, 43, 32, 23] show that selecting informative examples under limited compute can significantly improve generalization. In diffusion models, explicit data-centric work is more limited. D2C [19] condenses training sets via difficulty scoring; FiFA [45] filters preference data based on informativeness; and Alchemist [37] upgrades noisy public text-to-image data through iterative filtering. Mild corruption applied to pre-training data has also been shown to improve robustness [5]. Works like Ambient Diffusion [10] further demonstrate training on uncurated, bad data yields robust models but limited in image generation. For video generation, recent systems [39, 22] adopt multi-stage filtering pipelines evaluating visual quality, motion consistency, and caption alignment. Yet these pipelines discard videos whenever either VQ or MQ is low, meaning that samples strong in only one dimension are never exploited. In contrast, our approach explicitly leverages such imbalanced data by routing high-MQ/low-VQ and high-VQ/low-MQ videos to different regions of the denoising process, reducing reliance on rare jointly high-quality video samples.
@@ 2.3 Timestep and Loss Reweighting
Diffusion models learn across a sequence of noise levels, and many works redesign timestep sampling or loss weighting to improve training. Foundational improvements such as cosine schedules [31] and SNR-based loss reweighting [15] highlight that different timesteps contribute unequally to optimization. Perception Prioritized training [8] further adjusts weights to favor perceptually important denoising stages, while curriculum-based strategies [21] gradually expand the range of timesteps used during training to improve stability. However, these approaches apply a uniform timetable to all training samples: each image or video receives the same per-timestep weighting regardless of its content. Existing methods therefore optimize when the model should learn, but not what type of data should be emphasized at each timestep. Our TQD method differs by conditioning timestep allocation on sample-specific quality indicators, exploiting the natural hierarchy that early timesteps primarily shape motion while later timesteps refine appearance. This enables effective learning from imperfect data—something existing timestep curricula cannot address.

### 01 Apr 2026 v2

@@ 1 Introduction
In this work, we take a different perspective—shifting the question from which videos should be kept to how can we use imperfect videos more effectively? We begin by analyzing the learning behavior of video diffusion models and make two key observations. First, the denoising process exhibits hierarchical learning: motion emerges at high-noise timesteps while details refine at low-noise timesteps. Second, and more critically, through gradient-based analysis we discover that quality-imbalanced samples can match golden data’s gradients at appropriate timesteps. This reveals that imbalanced data need not be discarded but should be strategically allocated to matching denoising phases where they provide equivalent learning signals. Building upon this observation, we adopt timestep selection in the training process and instantiate it through Timestep-aware Quality Decoupling. Instead of sampling timesteps equally across all data, we adjust the sampling distribution based on each video’s quality profile. High-MQ videos are biased toward higher-noise timesteps, whereas high-VQ videos are biased toward lower-noise timesteps. This design allows the model to leverage imperfect data effectively, aligning each sample with the phase of diffusion training where it provides the most benefit.
@@ 2.2 Data Selection and Curation
Data curation plays an important role in large-scale generative modeling. In vision-language modeling, both curated datasets [35] and recent model-driven selection approaches [11, 43, 32, 23] show that selecting informative examples under limited compute can significantly improve generalization. In diffusion models, explicit data-centric work is more limited. D2C [19] condenses training sets via difficulty scoring; FiFA [45] filters preference data based on informativeness; and Alchemist [37] upgrades noisy public text-to-image data through iterative filtering. Mild corruption applied to pre-training data has also been shown to improve robustness [5]. Works like Ambient Diffusion [10] demonstrate that training on uncurated or imperfect data can yield robust models, highlighting the potential of data-centric approaches to handle noisy data. For video generation, recent systems [39, 22] adopt multi-stage filtering pipelines evaluating visual quality, motion consistency, and caption alignment. Yet these pipelines discard videos whenever either VQ or MQ is low, meaning that samples strong in only one dimension are never exploited. In contrast, our approach explicitly leverages such imbalanced data by routing high-MQ/low-VQ and high-VQ/low-MQ videos to different regions of the denoising process, reducing reliance on rare jointly high-quality video samples.
@@ 2.3 Timestep and Loss Reweighting
Diffusion models learn across a sequence of noise levels, and many works redesign timestep sampling or loss weighting to improve training. Foundational improvements such as cosine schedules [31] and SNR-based loss reweighting [15] highlight that different timesteps contribute unequally to optimization. Perception Prioritized training [8] further adjusts weights to favor perceptually important denoising stages, while curriculum-based strategies [21] gradually expand the range of timesteps used during training to improve stability. Most of these methods apply a global strategy shared across samples, without explicitly adapting timestep allocation to the characteristics of individual training examples.
Recent work has also explored timestep-dependent data utilization beyond such global designs. In particular, Ambient Diffusion Omni [10] demonstrates that selectively utilizing imperfect or out-of-distribution data across diffusion timesteps can improve optimization, highlighting the importance of timestep selection during training. Inspired by this line of research, we introduce timestep-dependent perspective into video training. Different from prior work that mainly studies general imperfect-data utilization, we identify the motion–visual quality dilemma in large-scale video data and propose Timestep-aware Quality Decoupling (TQD), which assigns sample-dependent timestep distributions according to motion and visual quality, enabling imbalanced video data to contribute more effectively during training.